# Structured coalescents, coagulation equations and multi-type branching processes

### v2

@@ Abstract.
Consider a structured population consisting of $d$ colonies, with migration rates that are proportional to a positive parameter $K$. We sample $N_{K}$ individuals distributed evenly across the $d$ colonies and trace their ancestral lineages back. Within a colony, we assume that pairs of ancestral lineages coalesce at a constant rate, as in a Kingman’s coalescent. We identify each ancestral lineage with the set, or block, of its descendants in the sample and we encode the state of the system using a $d$-dimensional vector of empirical measures; the $i$-th component records the blocks present in colony $i$ and the initial location of the lineages composing each block. We are interested in the asymptotic behaviour of the process of empirical measures as $K\to\infty$. We consider two regimes: the critical sampling regime, where $N_{K}\sim K$ and the large sampling regime where $N_{K}\gg K$. After an appropriate time-space scaling, we show that the process of empirical measures converges to the solution of a $d$-dimensional coagulation equation. In the critical sampling regime, the solution can be represented in terms of a multi-type branching process. In the large sampling regime, the solution can be represented in terms of the entrance law of a multi-type continuous state branching process.
@@ 1. Introduction
Coalescents, branching processes, and coagulation equations represent three fundamental approaches for modeling the dynamics of general interacting particle systems. Each of these captures a distinct yet interconnected facet of stochastic evolution: coalescents describe the merging of ancestral lineages backward in time [25, 24, 31, 38, 40]; branching processes model population growth and reproduction mechanisms forward in time [19, 3, 20, 27, 29, 15]; and coagulation equations provide deterministic approximations for the evolution of cluster sizes in systems undergoing mass-conserving mergers [26, 13, 12, 35].
The relationship between these objects is theoretically rich and practically significant. For example, it is well known [2, 11, 7] that the mean field limit of the, properly scaled, evolution of block sizes in Kingman’s coalescent is given by a Smoluchowski’s coagulation. Furthermore, probabilistic representations of one-dimensional coagulation equations and single-type branching processes can be found, for example, in [11]. In the multi-type setting these connections are more involved. For instance, in the recent works [22, 17] it has been shown that the genealogies of multi-type continuous state branching processes can be described (at least locally) by means of exchangeable multi-type coalescents [23, 32, 14]. Along the same lines, one of the present authors (E.S.) introduced a nested coalescent model [28, 9], where gene lineages evolve within a larger species tree. That study established a connection between the nested coalescent and a transport-coagulation equation, and demonstrated that the corresponding deterministic PDE admits a stochastic representation in terms of a branching CSBP. In this work, we establish a similar connection between a structured Kingman coalescent [36], a multidimensional Smoluchowski coagulation equation, and multi-type branching processes. This generalizes some of the aforementioned results available for the Kingman coalescent and the nested coalescent.
@@
A key assumption in the previous averaging result is that the sample size remains fixed as the migration scale $K$ increases. In contrast, this article focuses on a different asymptotic regime, where the sample size $N_{K}$ grows with the migration scale, i.e., $N_{K}\to\infty$ as $K\to\infty$. More precisely, we investigate two distinct sampling regimes: the critical sampling regime, where $N_{K}\sim K$, and the large sampling regime, where $N_{K}\gg K$. In both cases, each ancestral lineage retains information about the block of sampled individuals it traces back to and the aforementioned averaging does not hold anymore. To capture this result, we represent the system’s state as a $d$-dimensional vector of empirical measures, where the $i$-th component captures the blocks present in colony $i$ and their corresponding block configurations—the initial locations of the lineages within each block. We will show that this process converges to the solution of a multi-dimensional Smoluchowski-type coagulation equation under an appropriate small time scaling. Further, we will show that the resulting equation admits a natural probabilistic interpretation in terms of multi-type branching processes. Taken together, our approach hints at a unified framework that connects structured coalescents, coagulation equations, and multi-type branching processes.
@@ 2.1. The model
We consider a prototype model for the genealogies of a large population divided into $d$ colonies, where $d\in{\mathbb{N}}$ will remain fixed throughout the manuscript. The model takes the form of a structured coalescent that evolves according to the following principles. Within a colony, blocks coalesce as in Kingman’s coalescent (i.e. at a constant rate per pair of blocks), while blocks can also migrate between colonies at rates proportional to a scaling parameter $K$. The aim of this article is to analyze the effect of sample size $N_{K}$ in the corresponding ancestral structures, at small times, as $K\to\infty$.
Let us now formalize the previous description. Each individual in the population is identified by a unique number in $[N_{K}]\coloneqq\{1,\ldots,N_{K}\}$; each colony is identified by a unique number in $[d]$, which we will often refer to as a color. The state of the system is encoded by a structured partition as defined below.
Definition 2.1 (Structured/colored partition).
Let $N\in\mathbb{N}$. We refer to the elements of ×[d]$\mathbb{N}\times[d]$ as colored (positive) integers. We say that $S\subset[N]\times[d]$ is a coloring of $[N]$ if $S=\{(1,c_{1}),\ldots,(N,c_{N})\}$ for some $c_{1},\ldots,c_{N}\in[d]$. We refer to ${\boldsymbol{\pi}}\coloneqq(\pi_{i})_{i\in[d]}$ as a colored partition of a coloring $S$ of $[N]$ if $\pi_{1},\ldots,\pi_{d}$ are disjoint (possibly empty) collections of non-empty subsets of $S$ and $\pi_{1}\cup\pi_{2}\cup\ldots\cup\pi_{d}$ is a partition of $S$; we call the elements of $\pi_{i}$ blocks of color $i$. If $S$ is a coloring of $[N]$, we denote by ${\mathcal{S}}_{S}^{N}$ the set of all colored partitions of $S$ and by ${\mathcal{S}}^{N}$ the set of all colored partitions of some coloring of $[N]$. We equip ${\mathcal{S}}_{S}^{N}$ and ${\mathcal{S}}^{N}$ with the discrete topology.
@@
$S={\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2},{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3,4},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}$
is a coloring of $[N=5]$ with $d=2$ colors.
${{\boldsymbol{\pi}}}=\underbrace{{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}}_{\pi_{1}}\underbrace{{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}4}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}}_{\pi_{2}}$
is a colored partition of $S$ with $2$ blue blocks and 1 green block.
Now, we describe the anticipated multi-type coalescent as a Markov process on the set of colored partitions. Fix $K\geq 1$ and $N_{K}\in\mathbb{N}$. Let ${\boldsymbol{W}}\coloneqq(w_{i,j})_{i,j\in[d]}$ be a primitive matrix and ${\boldsymbol{\alpha}}\coloneqq(\alpha_{i})_{i\in[d]}\in{\mathbb{R}}_{+}^{d}$. The multi-type coalescent ${\boldsymbol{\Pi}}^{K}\coloneqq(\mathbf{\Pi}^{K}(t))_{t\geq 0}$ is the continuous-time Markov chain on ${\mathcal{S}}^{N_{K}}$ which is characterized by the following dynamics:
@@
    At time 00, start with a coloring of $[N_{K}]$. The partition $\Pi_{0}$ is the set of singletons whose coloring coincides with the coloring of $[N_{K}]$. For instance, for the coloring of $S$ in Example 2.2,
    $\Pi^{K}_{0}={\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}4}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}.$
@@
    {15}{15}at rate ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}\to{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}\ \ \mbox{at rate $Kw_{1,2}$.}$
@@
If we envision the coalescent as a random ultrametric tree, the color of a block and its internal coloring can be interpreted as follows. The color of a block at time $t$ is the position (colony) of the corresponding lineage. The internal coloring provides the labels and colors of the leaves supported by this lineage $t$ units of time in the past.
@@
Since the matrix ${\boldsymbol{W}}$ is primitive, it follows that it admits a unique stationary distribution, which we denote by ${\boldsymbol{\xi}}\coloneqq(\xi_{i})_{i\in[d]}$.
@@ 2.2. The empirical measure
In this work we will focus on the asymptotic properties of a functional of the coalescent $\boldsymbol{\Pi}^{K}$, which will code for the colony sizes, but also for the color-configurations within blocks (see Fig. 1). To encode this information, we will need a few definitions. We will say that the configuration of a block is ${{\boldsymbol{k}}}=(k_{i})_{i\in[d]}\in[N_{K}]_{0}^{d}$ if and only if the block contains $k_{1}$ elements of color $1$, $k_{2}$ elements of color $2$ etc. For instance, the configuration of the blue block ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}3}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}$ is ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}(1},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4})}$.
Figure 1. An illustration of the multi-type coalescent for $d=3$, $N_{K}=10$. Blocks are classified according to their colors, which code for the different colonies. A change in color represents a migration event. Specifically, multi-colored squares indicate a migration from colony $i$ to colony $j$, where the upper color corresponds to the origin colony $i$, and the lower color to the destination colony $j$.
@@
$\ \nu_{i}(t,\{{\boldsymbol{k}}\})\ \coloneqq\ \#\{\mbox{blocks of color $i$ with configuration ${{\boldsymbol{k}}}$}\},\quad{{\boldsymbol{k}}}=(k_{i})_{i\in[d]}\in[N_{K}]_{0}^{d}.$
In words, $\nu_{i}(t,\{{\boldsymbol{k}}\})$ is the number of lineages at time $t$ located in colony $i$ carrying $k_{1}$ leaves of color $1$ etc. As anticipated at the beginning of the section, our aim is to understand the interplay between the sampling size $N_{K}$ and the scale $K$ at which migrations (change of color) occur in the model. In our analysis, we will therefore distinguish between the two regimes depending on the asymptotic behavior of $\gamma_{K}\coloneqq N_{K}/K$. More precisely, we consider
@@
In the critical sampling regime, the rates of migration and coalescence events will both be of the same order as long as the total number of blocks remains of order $\mathcal{O}(K)$. In contrast, in the large sampling regime, coalescence events will dominate the dynamics as long as the total number of blocks is of a higher order than $K$, and only when the total number of blocks reaches order $K$ will coalescence events and migrations synchronize on the same time scale. We will therefore slow down time by $1/K$, so that the time frame in which coalescence dominates the dynamics is very short, and hence the strengths of migration and coalescence equilibrate very quickly (see Fig. 2 for an illustration of the time scaling). As a consequence, in the large sampling regime, blocks will quickly become very large, and we will therefore need to introduce an appropriate space scaling. More precisely, and to avoid unnecessary case distinctions when scaling blocks, we introduce the space-scaling parameter
@@
With this intuition in mind, we consider the rescaled measure valued process $(\mu^{K}_{i}(t,{\rm d}x))_{t\geq 0}$ valued in ${\mathcal{M}}_{f}(\mathbb{R}_{+}^{d})$ equipped with weak topology as
$\ \ \ \mu_{i}^{K}(t,{\rm d}x)\ \coloneqq\frac{1}{K}\ \nu^{K}_{i}\left(\frac{t}{K},s_{K}\,{\rm d}x\right),\qquad i\in[d],t\geq 0.$

### v3

@@ Abstract.
Consider a structured population consisting of $d$ colonies, with migration rates proportional to a positive parameter $K$. We sample $N_{K}$ individuals, distributed evenly across the $d$ colonies, and trace their ancestral lineages backward in time. Within each colony, we assume that any pair of ancestral lineages coalesces at a constant rate, as in Kingman’s coalescent. We identify each ancestral lineage with the set, or block, of its sampled descendants, and we encode the state of the system using a $d$-dimensional vector of empirical measures; the $i$-th component records the blocks present in colony $i$ together with the initial locations of the lineages composing each block.
We are interested in the asymptotic behavior of the process of empirical measures such as $K\to\infty$. We consider two regimes: the critical sampling regime, where $N_{K}\sim K$, and the large-sample regime, where $N_{K}\gg K$. After an appropriate time rescaling, we show that the process of empirical measures converges to the solution of a $d$-dimensional coagulation equation. In the critical sampling regime, the solution can be represented in terms of a multi-type branching process. In the large-sample regime, the solution can be represented in terms of the entrance law of a multi-type continuous-state branching process.
@@ 1. Introduction
Coalescents, branching processes, and coagulation equations represent three fundamental approaches to modeling the dynamics of interacting particle systems. Each captures a distinct yet interconnected facet of stochastic evolution: coalescents describe the merging of ancestral lineages backward in time [27, 26, 33, 40, 42]; branching processes model population growth and reproduction mechanisms forward in time [20, 3, 21, 29, 31, 16]; and coagulation equations provide deterministic approximations for the evolution of cluster sizes in systems undergoing mass-conserving mergers [28, 13, 12, 37].
The relationships between these objects are theoretically rich and practically significant. For example, it is well known [2, 11, 7] that the mean-field limit of the properly scaled evolution of block sizes in Kingman’s coalescent is given by a Smoluchowski coagulation equation. Furthermore, probabilistic representations of one-dimensional coagulation equations in terms of single-type branching processes can be found, for instance, in [11]. In the multi-type setting, these connections are more involved. For example, recent works [23, 18] show that the genealogies of multi-type continuous-state branching processes can be described (at least locally) by exchangeable multi-type coalescents [24, 34, 15]. Along similar lines, one of the present authors (E.S.) introduced a nested coalescent model [30, 9], in which gene lineages evolve within a larger species tree. That work established a connection between the nested coalescent and a transport–coagulation equation, and demonstrated that the corresponding deterministic PDE admits a stochastic representation in terms of a branching CSBP.
In this work, we establish a related connection between a structured Kingman coalescent [38], a multidimensional Smoluchowski coagulation equation, and multi-type branching processes. Our results generalize some of the aforementioned results available for the Kingman coalescent and the nested coalescent to a multi-type (or structured) setting, and allow for sample-size scaling regimes that differ from those studied in the classical literature.
@@
A key assumption in the previous averaging result is that the sample size remains fixed as the migration scale $K$ increases. In contrast, this article focuses on a different asymptotic regime, where the sample size $N_{K}$ grows with the migration scale, i.e., $N_{K}\to\infty$ as $K\to\infty$. More precisely, we investigate two distinct sampling regimes: the critical sampling regime, where $N_{K}\sim K$, and the large sampling regime, where $N_{K}\gg K$. In both cases, each ancestral lineage retains information about the block of sampled individuals it traces back to and the aforementioned averaging does not hold anymore. To capture this result, we represent the system’s state as a $d$-dimensional vector of empirical measures, where the $i$-th component captures the blocks present in colony $i$ and their corresponding block configurations – the initial locations of the lineages within each block. We will show that this process converges to the solution of a multi-dimensional Smoluchowski-type coagulation equation under an appropriate small time scaling. Further, we will show that the resulting equation admits a natural probabilistic interpretation in terms of multi-type branching processes. Taken together, our approach hints at a unified framework that connects structured coalescents, coagulation equations, and multi-type branching processes.
@@ 2.1. The model
We consider a structured coalescent with $d$ colonies, where $d\in\mathbb{N}$ is fixed throughout the manuscript. The process evolves as follows. Within each colony, blocks coalesce as in Kingman’s coalescent (i.e. at a constant rate per pair), and blocks migrate between colonies at rates proportional to a scaling parameter $K$. As shown, for example, in [35], such models arise naturally as genealogies of population-level systems. Although we use the terminology of “colonies”, the same model also describes a multi-type population, with types playing the role of colonies; migration then corresponds to mutation, and coalescence occurs within types exactly as within colonies. The aim of this article is to analyze how the sample size $N_{K}$ affects the corresponding ancestral structures at small times as $K\to\infty$.
Let us now formalize the previous description. Each individual in the population is identified by a unique number in $N_{K}]\coloneqq\{1,\ldots,N_{K}\}$; each colony is identified by a unique number in $[d]$, which we will often refer to as a color. The state of the system is encoded by a colored partition as defined below.
Definition 2.1 (Colored partition).
Let $N\in\mathbb{N}$. We refer to sets of the form $S=\{(1,c_{1}),\ldots,(N,c_{N})\}$, $c_{1},\ldots,c_{N}\in[d]$, as a coloring of $[N]$. We call $\boldsymbol{\pi}\coloneqq(\pi_{i})_{i\in[d]}$ a colored partition of a coloring $S$ of $[N]$ if $\pi_{1},\ldots,\pi_{d}$ are disjoint (possibly empty) collections of non-empty subsets of $S$ and if $\pi_{1}\cup\pi_{2}\cup\cdots\cup\pi_{d}$ is a partition of $S$; the elements of $\pi_{i}$ are called blocks of color $i$. If $S$ is a coloring of $[N]$, we denote by $\mathcal{S}_{S}^{N}$ the set of all colored partitions of $S$, and by $\mathcal{S}^{N}$ the set of all colored partitions of some coloring of $[N]$. We equip both $\mathcal{S}_{S}^{N}$ and $\mathcal{S}^{N}$ with the discrete topology.
@@
The set
[…]
is a coloring of $[5]$ with $d=2$ colors, with the coloring represented directly in scriptcolor rather than by adding an additional coordinate encoding the color; blue represents colour 1 and green represents colour 2.
[…]
is a colored partition of $S$ with $2$ blue blocks and 1 green block. Note that the coloring operates on two levels: the individual elements within each block are colored, and the blocks themselves receive their own colors.
We now describe the structured coalescent as a Markov process on the space of colored partitions. Fix $K\geq 1$ and $N_{K}\in\mathbb{N}$. Let $\boldsymbol{W}\coloneqq(w_{i,j})_{i,j\in[d]}$ be a primitive matrix (that is, $\boldsymbol{W}$ has nonnegative entries and there exists $n\in\mathbb{N}$ such that $\boldsymbol{W}^{n}$ has all entries strictly positive), and let $\boldsymbol{\alpha}\coloneqq(\alpha_{i})_{i\in[d]}\in\mathbb{R}_{+}^{d}$. The structured coalescent $\boldsymbol{\Pi}^{K}\coloneqq(\boldsymbol{\Pi}^{K}(t))_{t\geq 0}$ is the continuous-time Markov chain on $\mathcal{S}^{N_{K}}$ defined by the following dynamics:
@@
    At time 0, start with a coloring of $N_{K}]$. The partition $\Pi_{0}$ is the set of singletons whose coloring coincides with the coloring of $N_{K}]$. For instance, for the coloring of $S$ in Example 2.2,
[…]
@@
[…]
@@
If we envision the coalescent as a random ultrametric tree, the color of a block and its internal coloring can be interpreted as follows. The color of a block at time $t$ is the position (colony) of the corresponding lineage. The internal coloring records the labels and colors of the leaves supported by this lineage at time $t$ in the past.
@@
Since the matrix $\boldsymbol{W}$} is primitive, it has a unique stationary distribution, denoted by ${\boldsymbol{\xi}}\coloneqq(\xi_{i})_{i\in[d]}$, characterized by
$\sum_{j\in[d]\setminus\{i\}}\xi_{j}w_{j,i}=\xi_{i}\sum_{j\in[d]\setminus\{i\}}w_{i,j}.$
@@ 2.2. The empirical measure
[…]
Figure 1. An illustration of the structured coalescent for $d=3$, $N_{K}=10$. Blocks are classified according to their colors, which code for the different colonies. A change in color represents a migration event. Specifically, multi-colored squares indicate a migration from colony $i$ to colony $j$, where the upper color corresponds to the origin colony $i$, and the lower color to the destination colony $j$.
@@
$\nu_{i}(t,\{{\boldsymbol{k}}\})\ \coloneqq\ \#\{\mbox{blocks of color $i$ with configuration ${{\boldsymbol{k}}}$}\},\quad{{\boldsymbol{k}}}=(k_{i})_{i\in[d]}\in[N_{K}]_{0}^{d}.\$
In words, $\nu_{i}(t,\{\boldsymbol{k}\})$ is the number of lineages at time $t$ located in colony $i$ that carry $k_{1}$ leaves of color $1$, $k_{2}$ leaves of color $2$, and so on. As anticipated at the beginning of this section, our aim is to understand the interplay between the sample size $N_{K}$ and the scale $K$ at which migrations (changes of color) occur in the model. We therefore distinguish between two regimes, depending on the asymptotic behaviour of $\gamma_{K}\coloneqq N_{K}/K$. More precisely, we consider
@@
In the critical sampling regime, migration and coalescence occur on the same time scale as long as the total number of blocks remains of order $\mathcal{O}(K)$. In the large-sampling regime, coalescence dominates as long as the number of blocks exceeds order $K$, and the two mechanisms act on a common time scale only once the number of blocks has fallen to order $K$. We therefore rescale time by a factor $1/K$. This rescaling allows us to focus on the phase in which coalescence and migration operate on comparable time scales, after a very brief initial transient during which coalescence prevails (see Fig. 2 for an illustration of the time scaling). As a consequence, in the large-sampling regime, blocks quickly become very large, and we therefore need to introduce an appropriate scaling of the blocks. More precisely, and to avoid unnecessary case distinctions when scaling blocks, we introduce the scaling parameter
@@
With this intuition in mind, we consider the rescaled process $(\mu^{K}_{i}(t,{\rm d}x))_{t\geq 0}$ valued in ${\mathcal{M}}_{f}(\mathbb{R}_{+}^{d})$ equipped with weak topology as
$\mu_{i}^{K}(t,{\rm d}x)\ \coloneqq\frac{1}{K}\ \nu^{K}_{i}\left(\frac{t}{K},s_{K}\,{\rm d}x\right),\qquad i\in[d],\ t\geq 0,$
where the measure $\nu_{i}^{K}\!\left(u,s_{K}\,{\rm d}x\right)$ denotes the pushforward of $\nu_{i}^{K}(u,{\rm d}x)$ under the map $x\mapsto s_{K}x$. We will often write $\boldsymbol{\mu}^{K}$ for the vector of empirical-measure processes $(\mu_{i}^{K})_{i\in[d]}$. It is important to note that three distinct scalings are involved: time and the total mass of the measure are each scaled by $1/K$, while the block configurations are scaled by $s_{K}$.
@@
2.3. The one-dimensional case (the critical sampling regime)
To illustrate the type of asymptotic behavior one may expect for $(\boldsymbol{\mu}^{K}(t,{\rm d}x))_{t\geq 0}$, we first recall known results in the critical sampling regime for $d=1$, before turning to our own contributions. Throughout the remainder of this section, we therefore assume that $d=1$ and $N_{K}=K$.
With only a single color, the coalescent no longer needs to be defined on colored partitions; in this case, $\boldsymbol{\Pi}^{K}=\Pi^{K}$ reduces to the classical Kingman coalescent with coalescence rate $\rho=\alpha_{1}$. Describing block configurations then amounts to specifying block sizes, and we write $\mu^{K}$ and $\nu^{K}$ for the analogues of $\boldsymbol{\mu}^{K}$ and $\boldsymbol{\nu}^{K}$ under this identification. In this setting, it follows from [7, 37] that the rescaled process $(\mu^{K}(t))_{t\geq 0}$ converges to the solution of the discrete coagulation equation
$\partial_{t}u(t,n)$ $=\rho\Big(\frac{1}{2}\,u\star u(t,n)-\langle u(t),1\rangle\,u(t,n)\Big),$
$u(0,n)$ $=\delta_{1,n},$
for all $t\geq 0$ and $n\in\mathbb{N}$. Furthermore, it is well known (see, for example, [11, 2]) that the solution of this equation admits a natural probabilistic interpretation in terms of a branching process.
To see this, define, for $\lambda\in[0,1)$,
$v(t,\lambda):=1-\langle u(t),1\rangle+\sum_{n\in\mathbb{N}}u(t,n)\lambda^{n}.$
The function $v(t,\cdot)$ is the probability generating function of the probability distribution
$\big(1-\langle u(t),1\rangle,\ (u(t,n))_{n\in\mathbb{N}}\big)$
on $\mathbb{N}_{0}$, and clearly $v(0,\lambda)=\lambda$. A direct computation shows that, for all $t\geq 0$,
$\partial_{t}v(t,\lambda)$ $=\rho\left(\frac{1}{2}v(t,\lambda)^{2}+\frac{1}{2}-v(t,\lambda)\right)=h\big(v(t,\lambda)\big)-\rho\,v(t,\lambda),$
where $h(x)=\frac{\rho}{2}(x^{2}+1)$ is the probability generating function of $\frac{\rho}{2}\delta_{0}+\frac{\rho}{2}\delta_{2}$.
Since $v(0,\lambda)=\lambda$, this evolution coincides with the generating function of a continuous-time branching process in which each individual either gives birth to two offspring or dies, each at rate $\rho/2$, started from a single ancestor (see the backward Kolmogorov equation, e.g. [3, Chap. III.3, Eq. (5)]). For such a process, it is well known (see, e.g. [25, Eq. (17)]) that the one-dimensional marginals are geometric. Specifically, for all $n\in\mathbb{N}$ and $t\geq 0$,
$u(t,n)=q_{t}^{2}(1-q_{t})^{\,n-1},$
where $q_{t}=1/(1+\rho t/2)$ is the survival probability at time $t$, and $(q_{t}(1-q_{t})^{n-1})_{n\in\mathbb{N}}$ is the distribution of the branching process at time $t$ conditioned on survival.
With this in mind, we are now prepared to state our main convergence results.